# TaskEval: Assessing Difficulty of Code Generation Tasks for Large Language Models

### v2

@@ 3.1. General Overview of TaskEval
Figure 1. Overview of TaskEval Framework.
A general overview of TaskEval is given in Figure 1. TaskEval is operated in four steps: 1 First, transformations on the prompts are identified. Those transformations are any dimension that can lead to a variety of prompts for an LLM while keeping the essence of the task. In a way, such transformations can be seen as some metamorphic transformations (Chen et al., 2018) as they should not alter the semantics of the task. In our study, we focus on two transformations: Rephrasing and Context Information. Those transformations are then used in conjunction with each task to have a Prompt LLM generate prompts following the transformations’ specifications. 2 Secondly, the obtained prompts are fed into multiple CodeLLMs to generate code. The goal is to evaluate how each CodeLLM performs when confronted with different prompts. Using multiple CodeLLMs allows us to capture the diversity of architectures and training settings, which may respond differently to a given prompt. From the generated code, various metrics are extracted to assess performance for each task in the benchmark 3a. More precisely, task prompts will be processed using a topic modeling algorithm to cluster them into broader categories (e.g., sorting algorithms, SQL queries, etc.). The generated code will be parsed using its Abstract Syntax Tree (AST), and we have collected human annotators’ difficulty assessments for comparison. More details can be found in Section 4.5. In parallel, various IRT models (Baker, 2001) are fitted on the data. These models estimate latent attributes of items and respondents based on observed responses. In our case, the tasks serve as items, while the CodeLLMs act as respondents 3b. Once we have collected both the parameters and artifacts of the tasks, we can perform various analyses. First, we examine the tasks in the HumanEval+ and ClassEval benchmarks based on their characteristics. This enables us to compare them and establish a mapping between the two benchmarks. Then, we use our collected metrics (e.g., topics, AST structures, and human annotations) to further describe and contrast tasks both within and across benchmarks 4.
@@ 3.2. Generating Different Prompts based on Transformations
First, we use the Context Information transformation. Indeed, a straightforward way to act on the variability of a prompt is to simply disclose or withdraw certain information in the prompt, while not changing the task. EvoEval (Xia et al., 2024), for instance, proposes also to act on the difficulty. However, to modify the difficulty, they do so by adding/removing constraints in the prompt, thus modifying the semantics of the task, which is not desirable in our case. In our study, we instead propose to divide the information provided in the prompt into three levels. Those three levels of prompts are generated incrementally using the previous level as the starting point. We assume that prompts with more information should be more likely to lead to correct code for CodeLLMs (Deng et al., 2023; Xia et al., 2024). The levels are defined such as:
  • First level: contains a minimal amount of information that is the inputs/outputs of the task as well as a high-level description of what is intended. We could use the original prompt in benchmarks for each task as the first level, yet Siddiq et al. (Siddiq et al., 2024) showed that there can be inconsistencies across tasks’ prompts in benchmarks. Thus, we prefer to generate a new prompt to have similar formatting across all tasks and levels to reduce biases. This level is the closest one to the original prompt in benchmarks and contains the same level of information regarding the tasks.
@@
  • Third level: complete the prompt with direct references to the function such as variable names, and helper functions used similarly to a pseudo-algorithm but in natural language.
We chose to devise these three levels based on our observations. They prove to generate a sufficient amount of fine-grainess and to discriminate between different levels of information. Trying to fit additional level (e.g, between level 1 and 2) did not lead to noticeable changes. We present an example of different prompt levels obtained for a task in Listing 3.
@@
The second transformation used is Rephrasing which is paraphrasing a given prompt. Indeed, how the prompts or instructions are formulated has been shown to have an impact on LLM responses (Mizrahi et al., 2024; Weber et al., 2023; Wang et al., 2022b). Thus, similarly to these studies, we try to see if different ways of rephrasing a given prompt for a CodeLLM can affect how the CodeLLM addresses the task. To rephrase our prompts, we will use a similar approach as done by Gonen et al. (Gonen et al., 2022), that is prompting an LLM to rephrase a given text. Other techniques could be used to assist in rephrasing such as Chains-Of-Thoughts (Wei et al., 2022) but assessing the effectiveness of different rephrasing techniques is out-of-scope of this study.
@@
Using those two transformations, we come up with general templates that can be used to generate prompts (see replication package (rep, 2024)). Such general templates are then tailored for each task in the benchmark under consideration. Those templates are fed to a Prompt LLM (see Figure 1) to generate different prompts for a given task. The way prompts are generated is as follows: we start from the original docstring to have a common starting point for all prompts. Then, we apply the template to obtain different prompt levels with Context Information for a given task. Then, we apply the obtained prompts to the Rephrasing template. In our case, we first remove the examples from the original prompt, if any, before using our templates with the Prompt LLM. While removing examples can decrease the performance of CodeLLMs (Liu et al., [n. d.]), this allows for better control over the level of information we inject in our prompts. Indeed, examples can vary from simple basic examples to showcasing corner cases of the task. As such, this might artificially bias the output of CodeLLMs and the potential difficulty of the task because of the quality of the examples.
@@
At each step of our generation, that is after using each template on the prompts, we do a sanity check by manually checking them. This is accomplished by the first author, and cross-checked by the second author independently. We make sure that the Context Information levels are respected and the Rephrasing does not alter the semantics of the tasks. This is done in order not to have tasks that are artificially considered hard further down the line because of misleading prompts for the CodeLLM and not because of the task itself.
@@ 3.3. Collecting metrics for code snippets
The second step is to gather metrics for a given task. To do so, we leverage several CodeLLMs that will serve as our evaluators for code generation. For each prompt per task, each CodeLLM is asked to generate code with varying seeds using the sampling mode of CodeLLMs. By seed, we mean a random number used in the sampling-based generation mode of CodeLLMs. The rationale is that most CodeLLMs, especially non-open-source ones, are used with this non-greedy approach in order to generate code. Moreover, using the sampling method for the generation allows us to have a wider variety of generated code in order to evaluate the difficulty of each task. Indeed, we assume an easier task should not be as impacted by the stochasticity of the generation compared to a harder task. The obtained code is then processed and executed. Sample codes were then executed against available test inputs in the benchmark.
Once we have obtained all code for all prompts of all tasks and executed them against tests, we collect different metrics. In our case, we are mainly interested in the Functional correctness of a code sample. Functional correctness quantifies if a given code sample is correct or not using available test cases. Formally, the Functional correctness $Func$ for a code sample $s$ generated by CodeLLM $LLM$:
@@
where $T$ is the set of test cases symbolized by an input $i$ and an output $o$. That is, the code sample passes all test cases, i.e., gets a value of $1$. Intuitively, we expect a harder task to lead to a lower number of prompts leading to correct codes across rephrasing and different levels of context information.
@@
where $N_{j}$ is the total number of prompts for the task $j$. The final score $score_{LLM}(s_{j})$ represents the score for the task $j$ using a given LLM. We will use the scores for all CodeLLMs to determine the parameters of the task $j$.
@@ 3.4. Item Response Theory modelling
Since we have a metric measured across all tasks and multiple CodeLLMs, we can use these values as observed responses in an IRT model. IRT models, commonly used for analyzing and scoring psychological tests (Zanon et al., 2016), have already been used in NLP settings (Rodriguez et al., 2021a; Lalor et al., 2016). In the literature, several IRT models exist depending on the observed response distribution or the number of parameters of the model. A widely used model is the 2-Parameter (2PL) IRT model. Consider $j=1,...,m$ items being assessed by $i=1,...,n$ respondents, for example, students answering questions in an exam, leading to a true or false outcome $x_{ij}={0,1}$, thus following a Bernoulli distribution of parameter $p_{ij}$. The 2PL IRT model in this case is defined as:
@@
where $\theta_{i}$ is the ability of respondent $i$, $\delta_{j}$ is the difficulty, and $a_{j}$ is discriminant of item $j$. The IRT model maps a respondent’s ability to an expected response using the observed response through a logistic function with a difficulty parameter $\delta_{j}$ that indicates the ”location” on the difficulty range, and a discriminant parameter $a_{j}$, discriminating items based on their difficulty and the ability of the respondents, reflected in the steepness of the slope of the item characteristic curves. Each parameter can be estimated using maximum likelihood. In our case, we do not have a binary response but a continuous one, as the aggregated score at the task level has a value in [0,1]. As such, we will rely on the $\beta^{3}$-IRT model proposed by Chen et al. (Chen et al., 2019) which models the observed response as a continuous value in $(0,1)$ using a Beta distribution. The expected response for the model is defined as:
@@
where $p_{ij}$ is the response of respondent $i$ for item $j$ defined as $p_{ij}\sim B(\frac{\theta_{i}}{\delta_{j}}^{a_{j}},\frac{1-\theta_{i}}{1-\delta_{j}}^{a_{j}})$, $\theta_{i}\sim Beta(1,1)$, $\delta_{j}\sim Beta(1,1)$ and $a_{j}\sim Normal(1,1)$. In our approach, a respondent will be one of our CodeLLMs, an item one task from a benchmark, and $p_{ij}$ is the score $score_{LLM}(s_{j})$ on this task by the given CodeLLM as calculated in Equation 2. The parameters are estimated based on the observed response of the CodeLLM (i.e., the score) using Maximum Likelihood Estimation.
To illustrate the response obtainable given the different parameters, in Figure 2, we present a theoretical Item Characteristic Curve (ICC) mapping the ability of a CodeLLM with ability $\theta_{i}$ to the expected response $\mathop{\mathbb{E}}[p_{ij}|\theta_{i},\delta_{j},a_{j}]$ following the Equation 4 for four different tasks (in plain traits in Figure 2, each with its values of $\delta_{j}$ and $a_{j}$). It should be noted that since we are using functional correctness as our metric, the expected response translates to the likelihood of generating a correct code for a given task and CodeLLM, which is easily interpretable.
Figure 2. ICC between ability $\theta_{i}$ of a CodeLLM and the expected response $\mathop{\mathbb{E}}[p_{ij}|\theta_{i},\delta_{j},a_{j}]$ for four tasks with different $\delta_{j}$ and $a_{j}$ values.
@@
For a task $j$, $\delta_{j}$ and $a_{j}$ will ultimately affect how likely a given CodeLLM is to answer this task according to the IRT model: for instance, still for the blue task, a CodeLLM of ability $0.7$ has an expected response of $0.7$. For tasks that have the same $a_{j}$ but different $\delta_{j}$, the expected response is always higher for a CodeLLM with higher ability $\theta_{i}$: for instance, the expected response on the green task is always higher compared to the black task no matter the CodeLLM ability, as the difficulty of the green task is lower. In that sense, the difficulty is enough to characterize a task. This is particularly used in the 1PL IRT model (i.e., Equation 4 with $\forall j,a_{j}=1$). However, while it is easier to analyze, it is not necessarily the most realistic modelling. For instance, some tasks can have a disproportionate higher/lower expected response for different CodeLLMs of different capacities. This can arise, for instance, if these CodeLLMs were less trained on data similar to this particular task, and are more sensitive to how the task could be phrased. To account for this fact, the 2PL model offers a discriminant coefficient $a_{j}$ for each task.
When the discriminant starts changing, a higher difficulty equaling a lower expected response no matter the CodeLLM is not always true: still in Figure 2, despite the black task being “easier” than the red one (difficulty of $0.5$ compared to $0.6$), CodeLLMs with an ability higher than $0.6$ will have a higher expected response on the red task. Effectively, the red task is more discriminant than the blue task, as it better separates CodeLLMs according to their capacity. As $a_{j}\in\mathbb{R}$, it is even possible for $a_{j}<0$ (green curve in Figure 2): in that case, a task can lead to a higher expected answer for a CodeLLM with a lower ability even with a higher difficulty. That last case can be indicative of potential noise (Chen et al., 2019), annotation errors (Rodriguez et al., 2021a) or could point towards actual tasks that less capable CodeLLMs have an easier time on for several possible reasons: more facilities with the given input prompts, non-determinism in the generation, example that was memorized or that is similar to what the model was trained on.

### v3

@@ 3.1. General Overview of TaskEval
Figure 1. Overview of TaskEval Framework.: TaskEval framework process
Data: Benchmark under-test $D$ with $n$ tasks; Set of prompts transformations $\{t_{1},...,t_{T}\}$; Prompt LLM $Prompt_{LLM}$; Set of Code LLMs $\{CodeLLM_{1},...,CodeLLM_{m}\}$; Number of temperature generation $S$
Result: IRT model parameters $\delta$, $\alpha$, $\theta$
Prompt Generation
Initialize promptsArray an empty 2D array of size $n\times T$
for i $\leftarrow$ 1 to $n$ do
 basePrompt $\leftarrow$ $D_{i}$
 for j $\leftarrow$ 1 to $T$ do
    basePrompt $\leftarrow$ $Prompt_{LLM}$(basePrompt, $t_{j}$)
    promptsArray[i][j] $\leftarrow$ basePrompt
/* Manual verification occurs here */
Code Generation
Initialize scoreArray, a zero-filled 2D array of size $m\times n$
for i $\leftarrow$ 1 to $m$ do
 for j $\leftarrow$ 1 to $n$ do
    for t $\leftarrow$ 1 to $T$ do
       for s $\leftarrow$ 1 to $S$ do
          codeGenerated $\leftarrow$ $CodeLLM_{i}$(promptsArray[j][t], s)
          /* Using test oracles from benchmark $D$ */
          scoreArray[i][j] = scoreArray[i][j] + isCodeCorrect(codeGenerated)
    scoreArray[i][j] = scoreArray[i][j] / ($T\times S$)
IRT Compute
/* Computing for $n$ items and $m$ respondents */
$\delta$, $\alpha$, $\theta$ $\leftarrow$ IRT(scoreArray)
Algorithm 1 TaskEval high-level algorithm
A general overview of TaskEval is given in Figure 1 and on the pseudo-algorithm 1. TaskEval is operated in three steps: 1 First, transformations on the prompts are identified. Those transformations are any dimension that can lead to a variety of prompts for an LLM while keeping the essence of the task. In a way, such transformations can be seen as some metamorphic transformations (Chen et al., 2018) as they should not alter the semantics of the task. In our study, we focus on two transformations: Rephrasing and Context Information. Those transformations are then used in conjunction with each task to have a Prompt LLM generate prompts following the transformations’ specifications. 2 Secondly, the obtained prompts are fed into multiple CodeLLMs to generate code. The goal is to evaluate how each CodeLLM performs when confronted with different prompts. Using multiple CodeLLMs allows us to capture the diversity of architectures and training settings, which may respond differently to a given prompt. Generated codes are evaluated on available tests, and their results are aggregated into a score per task and per CodeLLM. Using these scores, an Item Response Theory (Baker, 2001) (IRT) model is trained 3. The model estimates the latent attributes of items (tasks) and respondents (CodeLLMs) based on the observed responses of all tasks and all CodeLLMs. In the following, we describe in more detail the different parts of our approach.
@@ 3.2. Generating Different Prompts based on Transformations
First, we use the Context Information transformation. Indeed, a straightforward way to act on the variability of a prompt is to simply disclose or withdraw certain information in the prompt, while not changing the task. EvoEval (Xia et al., 2024), for instance, also proposes to act on the difficulty. However, to modify the difficulty, they do so by adding/removing constraints in the prompt, thus modifying the semantics of the task, which is not desirable in our case. In our study, we instead propose to divide the information provided in the prompt into three levels. Those three levels of prompts are generated incrementally using the previous level as the starting point. We assume that prompts with more information should be more likely to lead to correct code for CodeLLMs (Deng et al., 2023; Xia et al., 2024). The levels are defined as:
  • First level: contains a minimal amount of information that is the inputs/outputs of the task, as well as a high-level description of what is intended. We could use the original prompt in benchmarks for each task as the first level, yet Siddiq et al. (Siddiq et al., 2024) showed that there can be inconsistencies across tasks’ prompts in benchmarks. Thus, we prefer to generate a new prompt to have similar formatting across all tasks and levels to reduce biases. This level is the closest one to the original prompt in benchmarks and contains the same level of information regarding the tasks.
@@
  • Third level: complete the prompt with direct references to the function, such as variable names, and helper functions used similarly to a pseudo-algorithm but in natural language.
We chose to devise these three levels based on our observations. They prove to generate a sufficient amount of fine-grainedness and to discriminate between different levels of information. Trying to fit an additional level (e.g, between level 1 and 2) did not lead to noticeable changes. We present an example of different prompt levels obtained for a task in Listing 3.
@@
The second transformation used is Rephrasing, which is paraphrasing a given prompt. Indeed, how the prompts or instructions are formulated has been shown to have an impact on LLM responses (Mizrahi et al., 2024; Weber et al., 2023; Wang et al., 2022b). Thus, similarly to these studies, we try to see if different ways of rephrasing a given prompt for a CodeLLM can affect how the CodeLLM addresses the task. To rephrase our prompts, we will use a similar approach as done by Gonen et al. (Gonen et al., 2023), that is, prompting an LLM to rephrase a given text. More precisely, we prompt the PromptLLM with a general meta-prompt rephrasing instruction “You are given a Python function in ### Function and a prompt in ### Prompt. Your task will be to reword the given prompt in order to generate a variation of said prompt.”, alongside the actual oracle code of the task (in ### Function) and the prompt to be rephrased (in ### Prompt). The LLM is instructed to generate all rephrasings one after another, using one prompt. The full prompt used can be found in our replication package [8]. We give an example in Listing 4. Other techniques could be used to assist in rephrasing/information, such as Chains-Of-Thoughts (Wei et al., 2022), but assessing the effectiveness of additional techniques is out of the scope of this study.
1def unique(l: list):
2 Base: """ Write a function named ’unique’ that returns a sorted list of unique elements from a given list. """
3 Rephrasing 1: """ Develop a function titled ’unique’ which outputs a list sorted with distinct elements extracted from a provided list."""
4 Rephrasing 2: """ Construct a function called ’unique’ that delivers a sorted list containing only unique elements from an input list."""
5 Rephrasing 3: """ Create a function known as ’unique’ that yields a list, sorted and containing only unique items from a specified list."""
6 Rephrasing 4: """ Design the function ’unique’ to produce a list where elements are unique and sorted, derived from a given list."""
7 Rephrasing 5: """ Formulate a function named ’unique’ that generates a sorted list of distinct elements originating from an input list."""
8 return sorted(set(l))
Listing 4: Example of different prompt levels for the task unique from HumanEval+, with the oracle code at the bottom.
@@
Using those two transformations, we come up with general templates that can be used to generate prompts (see replication package (rep, 2024)). Such general templates are then tailored for each task in the benchmark under consideration. Those templates are fed to a Prompt LLM (see Figure 1) to generate different prompts for a given task. The way prompts are generated is as follows: we start from the original docstring to have a common starting point for all prompts. Then, we apply the template to obtain different prompt levels with Context Information for a given task. Then, we apply the obtained prompts to the Rephrasing template. In our case, we first remove the examples from the original prompt, if any, before using our templates with the Prompt LLM. While removing examples can decrease the performance of CodeLLMs (Liu et al., [n. d.]), this allows for better control over the level of information we inject into our prompts. Indeed, examples can vary from simple, basic examples to showcasing corner cases of the task. As such, this might artificially bias the output of CodeLLMs and the potential difficulty of the task because of the quality of the examples.
@@
At each step of our generation, that is, after using each template on the prompts, we do a sanity check by manually checking them. This is accomplished by the first author and cross-checked by the second author independently. We make sure that the Context Information levels are respected and the Rephrasing does not alter the semantics of the tasks. This is done in order not to have tasks that are artificially considered hard further down the line because of misleading prompts for the CodeLLM and not because of the task itself.
@@ 3.3. Collecting metrics for code snippets
The second step is to gather metrics for a given task. To do so, we leverage several CodeLLMs that will serve as our evaluators for code generation. For each prompt per task, each CodeLLM is asked to generate code with varying seeds using the sampling mode of CodeLLMs. By seed, we mean a random number used in the sampling-based generation mode of CodeLLMs. The rationale is that most CodeLLMs, especially non-open-source ones, are used with this non-greedy approach to generate code. Moreover, using the sampling method for the generation allows us to have a wider variety of generated code in order to evaluate the difficulty of each task. Indeed, we assume an easier task should not be as impacted by the stochasticity of the generation compared to a harder task. The obtained code is then processed and executed. Sample codes were then executed against available test inputs in the benchmark.
Once we have obtained all the codes for all prompts of all tasks and executed them against tests, we can collect different metrics. In our case, we are mainly interested in the Functional correctness of a code sample. Functional correctness quantifies whether a given code sample is correct or not using available test cases. Formally, the Functional correctness $Func$ for a code sample $s$ generated by CodeLLM $LLM$:
@@
where $T$ is the set of test cases symbolized by an input $i$ and an output $o$. That is, the code sample passes all test cases, i.e., gets a value of $1$. Intuitively, we expect a harder task to lead to fewer prompts producing correct codes across rephrasing and different levels of context information.
@@
where $N_{j}$ is the total number of prompts for the task $j$. The final score $score_{LLM}(s_{j})$ represents the score for the task $j$ using a given LLM. We will use the scores for all CodeLLMs to determine the task’s parameters $j$.
@@ 3.4. Item Response Theory modelling
Since we have a metric measured across all tasks and multiple CodeLLMs, we can use these values as observed responses in an IRT model. IRT models, commonly used for analysing and scoring psychological tests (Zanon et al., 2016), have already been used in NLP settings (Rodriguez et al., 2021a; Lalor et al., 2016). In the literature, several IRT models exist depending on the observed response distribution or the number of parameters of the model. A widely used model is the 2-Parameter (2PL) IRT model. Consider j=1,,mj=1,...,m items being assessed by i=1,,ni=1,...,n respondents, for example, students answering questions in an exam, leading to true or false outcome $x_{ij}=\{0,1\}$, thus following a Bernoulli distribution of parameter $p_{ij}$. The 2PL IRT model in this case is defined as:
@@
where $\theta_{i}$ is the ability of respondent $i$, $\delta_{j}$ is the difficulty, and $a_{j}$ is discriminant of item $j$. The IRT model maps a respondent’s ability to an expected response using the observed response through a logistic function with a difficulty parameter $\delta_{j}$ that indicates the ”location” on the difficulty range, and a discriminant parameter $a_{j}$, discriminating items based on their difficulty and the ability of the respondents, reflected in the steepness of the slope of the item characteristic curves. In our case, we do not have a binary response but a continuous one, as the aggregated score at the task level has a value in [0,1]. As such, we will rely on the $\beta^{3}$-IRT model proposed by Chen et al. (Chen et al., 2019), which models the observed response as a continuous value in $(0,1)$ using a Beta distribution. The expected response for the model is defined as:
@@
where $p_{ij}$ is the response of respondent $i$ for item $j$ defined as $p_{ij}\sim B(\frac{\theta_{i}}{\delta_{j}}^{a_{j}},\frac{1-\theta_{i}}{1-\delta_{j}}^{a_{j}})$, $\theta_{i}\sim Beta(1,1)$, $\delta_{j}\sim Beta(1,1)$ and $a_{j}\sim Normal(1,1)$. In our approach, a respondent will be one of our CodeLLMs, an item a task from a benchmark, and $p_{ij}$ is the score $score_{LLM}(s_{j})$ on this task by the given CodeLLM as calculated in Equation 2. The parameters are estimated based on the observed response of CodeLLM (that is, the score) using the maximum likelihood estimate.
To illustrate the response obtainable given the different parameters, in Figure 2, we present a theoretical Item Characteristic Curve (ICC) mapping the ability of a CodeLLM with ability $\theta_{i}$ to the expected response $\mathop{\mathbb{E}}[p_{ij}|\theta_{i},\delta_{j},a_{j}]$ following the Equation 4 for four different couple of parameters $(\delta_{j},a_{j})$ (in plain traits in Figure 2). The values are chosen to illustrate the effect of each parameter, as we detail in the following. It should be noted that since we are using functional correctness as our metric, the expected response translates to the likelihood of generating a correct code for a given task and CodeLLM, which is easily interpretable.
Figure 2. ICC between ability $\theta_{i}$ of a CodeLLM and the expected response $\mathop{\mathbb{E}}[p_{ij}|\theta_{i},\delta_{j},a_{j}]$ for four tasks with different $\delta_{j}$ and $a_{j}$ values.: Example of Item Characteristic Curves
@@
For a task $j$, $\delta_{j}$ and $a_{j}$ will ultimately affect how likely a given CodeLLM is to answer this task according to the IRT model: for instance, still for the blue task, a CodeLLM of ability $0.7$ has an expected response of $0.7$. For tasks that have the same $a_{j}$ but different $\delta_{j}$, the expected response is always higher for a CodeLLM with a higher ability $\theta_{i}$: for instance, the expected response on the green task is always higher compared to the black task, no matter the CodeLLM ability, as the difficulty of the green task is lower. In that sense, the difficulty is enough to characterise a task. This is particularly used in the 1PL IRT model (i.e., Equation 4 with $\forall j,a_{j}=1$). However, while it is easier to analyse, it is not necessarily the most realistic modelling. For instance, some tasks can have a disproportionately higher/lower expected response for different CodeLLMs of different capacities. This can arise, for instance, if these CodeLLMs were less trained on data similar to this particular task, and are more sensitive to how the task could be phrased. To account for this fact, the 2PL model offers a discriminant coefficient $a_{j}$ for each task.
When the discriminant starts changing, a higher difficulty equaling a lower expected response no matter the CodeLLM is not always true: still in Figure 2, despite the black task being “easier” than the red one (difficulty of $0.5$ compared to $0.6$), CodeLLMs with an ability higher than $0.6$ will have a higher expected response on the red task. Effectively, the red task is more discriminant than the black task, as it better separates CodeLLMs according to their capacity. As $a_{j}\in\mathbb{R}$, it is even possible for $a_{j}<0$ (green curve in Figure 2): in that case, a task can lead to a higher expected answer for a CodeLLM with a lower ability even with a higher difficulty. That last case can be indicative of potential noise (Chen et al., 2019), annotation errors (Rodriguez et al., 2021a) or could point towards actual tasks that less capable CodeLLMs have an easier time on for several possible reasons: more facilities with the given input prompts, non-determinism in the generation, example that was memorized or that is similar to what the model was trained on etc.